OpenStudy (anonymous):

you run 11s at a speed of 1m/s then run for 11s at 3.2m/s along a straight track what is the average velcity? help please

OpenStudy (tkhunny):

\(\dfrac{Total\;Distance}{Total\;Time}\) Calculate the two pieces and you'll be done.

OpenStudy (anonymous):

@tkhunny I'm not sure how to do that....I'm really confused

OpenStudy (tkhunny):

Distance = Rate * Time Put each circumstance in the formula. I'll do one. 11s at a speed of 1m/s Distance = 1 m/s * 11 s = 11 m Okay. Your turn. 11s at 3.2m/s

OpenStudy (anonymous):

3.2m/s * 11 =35.2

OpenStudy (tkhunny):

Okay. It should be "m", but that's good. Now, add up the total time and the total distance. You'll be done.